OpenStudy (anonymous):

y+x=-4 what is the y intercept

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

y=mx+b, b is the y-intercept, which is -4

OpenStudy (anonymous):

To find the y intercept you plug in a 0 for x. y + 0 = -4 y = -4

OpenStudy (anonymous):

how would u graph it. doesn't it need two integers to graph it??

OpenStudy (anonymous):

(x,y) = (0,-4) Remember that the x part is 0, and the y part is whatever you get when you plug in the 0 for x.

OpenStudy (anonymous):

ty. are u good at graphing??? I suck I have a few more problems can u help???